# Constraints from CMB lensing tomography with projected bispectra

### v1

@@ Abstract
We measure the angular power spectrum and bispectrum of the projected overdensity of photometric DESI luminous red galaxies, and its cross-correlation with maps of the Cosmic Microwave Background lensing convergence from Planck. This analysis is enabled by the use of the “filtered-squared bispectrum” approach, introduced in previous work, which we generalise here to the case of cross-correlations between multiple fields. The projected galaxy bispectrum is detected at very high significance (above $30\sigma$ in all redshift bins), and the galaxy-galaxy-convergence bispectrum is detected above $5\sigma$ in the three highest-redshift bins. We find that the bispectrum is reasonably well described over a broad range of scales by a tree-level prediction using the linear galaxy bias measured from the power spectrum. We carry out the first cosmological analysis combining projected power spectra and bispectra under a relatively simple model, and show that the galaxy bispectrum can be used in combination with the power spectrum to place a constraint on the amplitude of matter fluctuations, $\sigma_{8}$, an on the non-relativistic matter fraction $\Omega_{m}$. We find that data combinations involving the galaxy bispectrum recover constraints on these parameters that are in good agreement with those found from the traditional “2$\times$2-point” combination of galaxy-galaxy and galaxy-convergence power spectra, across all redshift bins.
@@ 1 Introduction
The analysis of projected probes of the large-scale structure (i.e. proxies of the matter fluctuations integrated along the line of sight and mapped onto the celestial sphere) has been one of the most productive areas of cosmology over the last decade. This is thanks to the wide variety of complementary probes we now have access to, and to the fast growth in the cosmic volume we can cover with them. Examples of these probes are the projected clustering of galaxies (Nicola et al., 2020; Vakili et al., 2023; Xu et al., 2023), the gravitational lensing of the Cosmic Microwave Background (CMB lensing) (Lewis & Challinor, 2006), galaxy weak lensing (Bartelmann & Schneider, 2001; Mandelbaum, 2018), or the Sunyaev-Zel’dovich effect (Carlstrom et al., 2002). The combination of galaxy clustering and CMB lensing has been particularly fruitful: by measuring the cross-power spectrum between maps of the CMB lensing convergence $\kappa$ and the overdensity of galaxies $\delta_{g}$ at different redshifts, and combining it with the auto-correlation of these galaxies, we can simultaneously determine the galaxy bias relation and measure the amplitude of matter fluctuations at the redshifts of these galaxies. This technique, often called “lensing tomography” or “$2\times 2$-point analysis” [^1] This nomenclature is imprecise and may lead to confusion. We only reluctantly subscribe to it due to its now prevalent use by the community. For clarity, in this work, when discussing “$2\times 2$-point” analyses, we will refer to the combination of the galaxy-galaxy and the galaxy-convergence power spectra, and not any other pair of cross-correlations., has now been used to place constraints on the growth history with percent-level precision spanning the last 12 billion years ($z\lesssim 3$) (Krolewski et al., 2021; White et al., 2022; Alonso et al., 2023; Farren et al., 2024; Sailer et al., 2025b; Qu et al., 2025; Sailer et al., 2025a; de Belsunce et al., 2025). Including the CMB lensing auto-correlation [^2] Often referred to as “$3\times 2$-point”, not to be confused with the same technique when applied to galaxy weak lensing – even though often more than 3 correlations in total are actually used in both cases., one can then extend the growth reconstruction to redshifts above that probed by the highest-redshift galaxy sample used (Farren et al., 2025).
@@
In spite of this success, further information may still be gained from the joint analysis of galaxies and CMB lensing. As the Universe evolves, the matter overdensities grow via non-linear gravitational collapse, which generates non-Gaussian features in their spatial distribution. In this regime, the two-point correlators used in the standard $2\times 2$-point analysis fail to capture all of the information present in both $\delta_{g}$ and $\kappa$. To recover this information, various techniques have been proposed in the literature. On the one hand, field-level analyses constitute the manifestly-optimal approach to extract this information, although they are difficult to deploy in practice, due to their computational complexity, and the need to simulate intricate astrophysical and observational effects present in the data (Jasche & Wandelt, 2013; Lavaux et al., 2019; Porqueres et al., 2022, 2023). The alternative approach is to employ well-educated summary statistics, commonly designed by performing non-linear (and sometimes non-local) operation on the data before compressing it onto a small set of measurements (e.g. some form of correlator) (Obreschkow et al., 2013; Friedrich et al., 2018; Eickenberg et al., 2022; Massara et al., 2023; Grewal et al., 2022; Oliveira Franco et al., 2022; Régaldo-Saint Blancard et al., 2024; Yip et al., 2024; Cowell et al., 2024). Within this category, the bispectrum (i.e. the Fourier- or harmonic-space third-order correlator) stands out for several reasons. First, the bispectrum is the lowest-order non-Gaussian summary statistic, and thus should contain significantly more information than other correlators for mildly non-Gaussian fields. Secondly, it is a theoretically tractable observable, in the sense that predictions for it can be made from first principles (e.g. via perturbation theory), without the need for simulations that must encompass all physical effects present in the data (Scoccimarro et al., 1999; Bernardeau et al., 2002; Simpson et al., 2011).
@@
Incorporating the bispectrum in cosmological analyses can be a significant challenge, however, in comparison with the standard power-spectrum-based studies. First, optimal bispectrum estimators are often significantly more computationally challenging than their power spectrum versions (Philcox, 2021; Philcox & Flöss, 2024). Furthermore, unlike the fairly compact power spectrum, information in the bispectrum measurement is spread over a large number of “triangle configurations”. Thirdly, it is often difficult to derive robust estimates of the bispectrum uncertainties, and its covariance with the power spectrum (Biagetti et al., 2022). Simulations may be used in this case, but a large number of them is needed to ensure convergence over the large number of triangle configurations available. Finally, bispectrum predictions may be slow to produce from first principles, mostly due again to the large number of configurations. Partly because of these difficulties, the projected bispectrum has not been widely exploited in cosmological large-scale structure analyses. Although the bispectrum has been considered in the context of weak lensing (Cooray & Hu, 2001; Takada & Jain, 2004; Munshi et al., 2020a), most existing constraints are based on specific real-space two-point correlations (Gatti et al., 2022; Secco et al., 2022; Porth et al., 2024; Burger et al., 2024). One of the few examples present in the literature is the recent measurements of the projected galaxy-galaxy-$\kappa$ bispectrum presented in Farren et al. (2023).
@@
Recently, Harscouet et al. (2025) (H25 hereafter) introduced the filtered-squared bispectrum estimator (FSB), addressing some of the difficulties mentioned above, and significantly simplifying the analysis of projected bispectra. The FSB is similar in spirit to the skew-spectrum (Cooray, 2001; Calabrese et al., 2010; Smidt et al., 2010; Moradinezhad Dizgah et al., 2020; Munshi et al., 2020b) (i.e. the cross-correlation between a given field and its square), introducing a filtering operation that allows it to recover different bispectrum configurations. Most importantly, by employing power spectrum techniques, H25 demonstrated that the FSB covariance (including its covariance with the power spectrum) may be accurately estimated using fast analytical methods in a completely data-driven manner, solving one of the main hurdles of bispectrum analysis. In this paper, we exploit the FSB to measure the galaxy-galaxy-galaxy and galaxy-galaxy-$\kappa$ bispectrum in real data, in combination with the standard $2\times 2$-point data vector. We will use these measurements to carry out a first tentative cosmological analysis combining projected power spectra and bispectra, quantifying the ability of the bispectrum to reduce the final parameter uncertainties, and to test the robustness of cosmological constraints, through the consistency between different data combinations.
@@
The paper is structured as follows. Section 2 presents the FSB estimator, and the extension we introduce here to the case of multi-field cross-correlations. It also describes the model we will use to interpret our measurements, and our parameter inference pipeline. The galaxy samples and CMB lensing data we use in the analysis are described in Section 3. Our results, including an assessment of the bispectrum detections, the complementarity of different data combinations, and the cosmological constraints derived from the measurements, are presented in Section 4. Finally, Section 5 summarises our results and discusses potential avenues for future exploration.
@@ 2.1.1 The estimator
Let us consider a scalar field $a(\hat{\bf n})$ defined on the sphere, and a filter $W_{\ell}^{L}$ which only has significant support over a compact range of angular scales $\ell\in L$. Note that this choice of filter is not unique, and that we can perform the following steps for a set of filters defined over different multipole ranges – but for the sake of simplicity we focus for now on one filter $W_{\ell}^{L}$. Now let us construct a ”filtered-squared” version of $a$, by first filtering it in harmonic space so as to only keep the information coming from the scales selected by the $W_{\ell}^{L}$ filter:
@@
The FSB is the angular power spectrum of $(a_{L}^{2})(\hat{\bf n})$ with the original field $a(\hat{\bf n})$: in the simplest case (that of full-sky observations, with the power spectrum estimated at each integer $\ell$), this is simply
@@
In the presence of a mask $w_{a}(\hat{\bf n})$ associated with field $a(\hat{\bf n})$, and using bandpowers (i.e. bins of $\ell$) labelled by an index $b$, we replace this by the pseudo-$C_{\ell}$ (PCL) estimator, as implemented in the NaMaster code [^3] https://github.com/LSSTDESC/NaMaster (Alonso et al., 2019). This version of the FSB estimator accounts for mode-coupling effects introduced when taking the power spectrum of two fields partially defined on the sky. However, the back and forth between real and harmonic space during the early filtering step inevitably introduces some additional mode coupling which is not fully accounted for in the estimator; H25 verified that the estimator remained robust to those residual mode coupling effects, by measuring the (largely negligible) effect caused by masks of varied complexity on the FSB estimator.
@@ 2.1.3 FSB covariance
    Pushing the FSB reinterpretation of the bispectrum as a power spectrum between $a$ and $a_{L}^{2}$, we can use existing techniques (García-García et al., 2019) to estimate the analytical Gaussian covariance (i.e. the result of assuming that both $a$ and $a_{L}^{2}$ are Gaussian fields). As shown in H25, this recovers all purely diagonal terms, both disconnected and connected, contributing to the covariance.
@@ 2.2 Multi-field FSB
Instead, in this paper we focus on the multi-field version of the estimator. In particular, we will consider cross-correlating a filtered-squared field $a^{2}_{L}(\hat{\bf n})$ with another field $f(\hat{\bf n})$, effectively probing their cross-bispectrum $b^{aaf}_{\ell_{1}\ell_{2}\ell_{3}}$. This measurement requires very few changes to the estimator: our original expression for the single-field FSB (Eq. 2) becomes
@@ 2.3 Maps and power spectrum measurements
  • To determine the detectability of the projected bispectrum, we will measure the FSB in four filters below the $2N_{\rm side}$ threshold. These four consecutive top-hat filters each have width $\Delta\ell=126$, partitioning the multipole range [2, $\ell\in[2,\,506]$.
@@
This must be done for the filters and bandpower window functions corresponding to any field affected by the PWF (i.e. the galaxy overdensity, but not $\kappa$ in our case). Power spectra should be likewise corrected by the PWF where relevant. Note that any pure shot noise contribution (namely the power spectrum shot noise, $N_{\ell}^{gg}$, and the $ggg$ FSB noise term in Eq. 33) should not be corrected for the PWF, as white noise spectra remain constant regardless of the pixelization scheme [^4] This can be understood in terms of a purely Poisson process, in which the measurements on different pixels remain statistically independent regardless of the pixel size..
@@ 2.4.1 Galaxy clustering and CMB lensing
We will use a perturbative expansion in Eulerian space, up to second order, to connect the galaxy and matter overdensities (Fry & Gaztanaga, 1993; McDonald & Roy, 2009). Specifically, in Fourier space this reads
@@ 2.4.2 Power spectrum modelling
where $k_{\ell}(\chi)\equiv(\ell+1/2)/\chi$ is the 3D wavenumber associated with the angular scale $\ell$, and $q_{\alpha}(\chi)$ is the radial kernel of the $\alpha$ field. This relation is accurate in the Limber approximation (Limber, 1953; LoVerde & Afshordi, 2008), which holds for the wide radial kernels used in this work.
@@
where $P^{mm}(k,z)$ is the matter power spectrum.
@@ 2.4.3 Bispectrum Modelling
In this work we will focus on the galaxy-galaxy-galaxy and galaxy-galaxy-convergence bispectra ($ggg$ and $gg\kappa$ hereafter), and hence must model the 3D $ggg$ and $ggm$ bispectra. We do not include the convergence-convergence-galaxy ($\kappa\kappa g$) bispectrum in this work, which might require a more complex modelling of the lensing reconstruction noise and is expected to have a lower signal than the $ggg$ and $gg\kappa$ bispectra. As in the case of the power spectrum, we will do this at lowest order in perturbation theory (i.e. tree level). Using the perturbative bias expansion introduced in Section 2.4.1, the corresponding bispectra are given by (see Schmittfull et al., 2015; Desjacques et al., 2018; Cagliari et al., 2025):
@@
$P_{k}$ denotes the linear matter power spectrum. The cyclical permutations for the $ggg$ bispectrum are, explicitly, $\{\mathbf{k},\mathbf{k}^{\prime}\}$, $\{\mathbf{k}^{\prime},\mathbf{k}^{\prime\prime}\}$, and $\{\mathbf{k}^{\prime\prime},\mathbf{k}\}$. In turn, the permutations in the $gg\kappa$ case are the same for the linear term, but only $\{\mathbf{k}^{\prime},\mathbf{k}^{\prime\prime}\}$ and $\{\mathbf{k}^{\prime\prime},\mathbf{k}\}$ contribute for the quadratic and tidal terms. Once computed, the 3D bispectrum $B^{ggg}(\mathbf{k},\mathbf{k}^{\prime},\mathbf{k}^{\prime\prime})$ ($B^{ggm}$) can be inserted into Eq. 2.4.3 to recover the projected bispectrum $b^{ggg}_{\ell\ell^{\prime}\ell^{\prime\prime}}$ ($b^{gg\kappa}_{\ell\ell^{\prime}\ell^{\prime\prime}}$).

### v2

@@ Abstract
We measure the angular power spectrum and bispectrum of the projected overdensity of photometric DESI luminous red galaxies, and its cross-correlation with maps of the Cosmic Microwave Background lensing convergence $\kappa$ from Planck. This analysis is enabled by the use of the “filtered-squared bispectrum” approach, introduced in previous work, which we generalise here to the case of cross-correlations between multiple fields. The projected galaxy bispectrum is detected at very high significance (above $30\sigma$ in all redshift bins), and the galaxy-galaxy-convergence bispectrum is detected above $5\sigma$ in the three highest-redshift bins. We find that the bispectrum is reasonably well described over a broad range of scales by a tree-level prediction using the linear galaxy bias measured from the power spectrum. We carry out the first cosmological analysis combining projected power spectra ($gg$ and $g\kappa$) and bispectra (gggggg only) under a relatively simple model, and show that the galaxy bispectrum can be used in combination with the power spectrum to place a constraint on the amplitude of matter fluctuations, $\sigma_{8}$, an on the non-relativistic matter fraction $\Omega_{m}$. We find that data combinations involving the galaxy bispectrum recover constraints on these parameters that are in good agreement with those found from the traditional $\times2$-point” combination of galaxy-galaxy and galaxy-convergence power spectra, across all redshift bins.
@@ 1 Introduction
The analysis of projected probes of the large-scale structure (i.e. proxies of the matter fluctuations integrated along the line of sight and mapped onto the celestial sphere) has been one of the most productive areas of cosmology over the last decade. This is thanks to the wide variety of complementary probes we now have access to, and to the fast growth in the cosmic volume we can cover with them. Examples of these probes are the projected clustering of galaxies (Nicola et al., 2020; Vakili et al., 2023; Xu et al., 2023), the gravitational lensing of the Cosmic Microwave Background (CMB lensing) (Lewis and Challinor, 2006), galaxy weak lensing (Bartelmann and Schneider, 2001; Mandelbaum, 2018), or the Sunyaev-Zel’dovich effect (Carlstrom et al., 2002). The combination of galaxy clustering and CMB lensing has been particularly fruitful: by measuring the cross-power spectrum between maps of the CMB lensing convergence $\kappa$ and the overdensity of galaxies $\delta_{g}$ at different redshifts, and combining it with the auto-correlation of these galaxies, we can simultaneously determine the galaxy bias relation and measure the amplitude of matter fluctuations at the redshifts of these galaxies. This technique, often called “lensing tomography” or $2\times 2$-point analysis” [^1] This nomenclature is imprecise and may lead to confusion. We only reluctantly subscribe to it due to its now prevalent use by the community. For clarity, in this work, when discussing $2\times 2$-point” analyses, we will refer to the combination of the galaxy-galaxy and the galaxy-convergence power spectra, and not any other pair of cross-correlations., has now been used to place constraints on the growth history with percent-level precision spanning the last 12 billion years ($z\lesssim 3$) (Krolewski et al., 2021; White et al., 2022; Alonso et al., 2023; Farren et al., 2024; Sailer et al., 2025b; Qu et al., 2025; Sailer et al., 2025a; de Belsunce et al., 2025). Including the CMB lensing auto-correlation [^2] Often referred to as $3\times 2$-point”, not to be confused with the same technique when applied to galaxy weak lensing – even though often more than 3 correlations in total are actually used in both cases., one can then extend the growth reconstruction to redshifts above that probed by the highest-redshift galaxy sample used (Farren et al., 2025).
@@
In spite of this success, further information may still be gained from the joint analysis of galaxies and CMB lensing. As the Universe evolves, the matter overdensities grow via non-linear gravitational collapse, which generates non-Gaussian features in their spatial distribution. In this regime, the two-point correlators used in the standard $2\times 2$-point analysis fail to capture all of the information present in both $\delta_{g}$ and $\kappa$. To recover this information, various techniques have been proposed in the literature. On the one hand, field-level analyses constitute the manifestly-optimal approach to extract this information, although they are difficult to deploy in practice, due to their computational complexity, and the need to simulate intricate astrophysical and observational effects present in the data (Jasche and Wandelt, 2013; Lavaux et al., 2019; Porqueres et al., 2022, 2023). The alternative approach is to employ well-educated summary statistics, commonly designed by performing non-linear (and sometimes non-local) operation on the data before compressing it onto a small set of measurements (e.g. some form of correlator) (Obreschkow et al., 2013; Friedrich et al., 2018; Eickenberg et al., 2022; Massara et al., 2023; Grewal et al., 2022; Oliveira Franco et al., 2022; Régaldo-Saint Blancard et al., 2024; Yip et al., 2024; Cowell et al., 2024). Within this category, the bispectrum (i.e. the Fourier- or harmonic-space third-order correlator) stands out for several reasons. First, the bispectrum is the lowest-order non-Gaussian summary statistic, and thus should contain significantly more information than other correlators for mildly non-Gaussian fields. Secondly, it is a theoretically tractable observable, in the sense that predictions for it can be made from first principles (e.g. via perturbation theory), without the need for simulations that must encompass all physical effects present in the data (Scoccimarro et al., 1999; Bernardeau et al., 2002; Simpson et al., 2011; Montandon et al., 2023).
@@
Incorporating the bispectrum in cosmological analyses can be a significant challenge, however, in comparison with the standard power-spectrum-based studies. First, optimal bispectrum estimators are often significantly more computationally challenging than their power spectrum versions (Scoccimarro, 2015; Philcox, 2021; Philcox and Flöss, 2024). Furthermore, unlike the fairly compact power spectrum, information in the bispectrum measurement is spread over a large number of “triangle configurations”. Thirdly, it is often difficult to derive robust estimates of the bispectrum uncertainties, and its covariance with the power spectrum (Biagetti et al., 2022). Simulations may be used in this case, but a large number of them is needed to ensure convergence over the large number of triangle configurations available. Similarly, the mode-coupling matrix – which accounts for mode loss and coupling induced by the survey geometry – is also numerically challenging to evaluate due to the sheer number of pairwise combinations of triangle configurations (although the survey window function has been studied for specific bispectrum statistics in e.g. Sugiyama et al., 2019). Finally, bispectrum predictions may be slow to produce from first principles, mostly due again to the large number of configurations. Partly because of these difficulties, the projected bispectrum has not been widely exploited in cosmological large-scale structure analyses. Although the bispectrum has been considered in the context of weak lensing (Cooray and Hu, 2001; Takada and Jain, 2004; Munshi et al., 2020b), most existing constraints are based on specific real-space two-point correlations (Gatti et al., 2022; Secco et al., 2022; Porth et al., 2024; Burger et al., 2024). One of the few examples present in the literature is the recent measurements of the projected galaxy-galaxy-$\kappa$ bispectrum presented in Farren et al. (2023).
@@
Recently, Harscouet et al. (2025) (H25 hereafter) introduced the filtered-squared bispectrum estimator (FSB), addressing some of the difficulties mentioned above, and significantly simplifying the analysis of projected bispectra. The FSB is similar in spirit to the skew-spectrum (Cooray, 2001; Calabrese et al., 2010; Smidt et al., 2010; Moradinezhad Dizgah et al., 2020; Munshi et al., 2020a) (i.e. the cross-correlation between a given field and its square), introducing a filtering operation that allows it to recover different bispectrum configurations. Most importantly, by employing power spectrum techniques, H25 demonstrated that the FSB covariance (including its covariance with the power spectrum) may be accurately estimated using fast analytical methods in a completely data-driven manner, solving one of the main hurdles of bispectrum analysis. In this paper, we exploit the FSB to measure the galaxy-galaxy-galaxy and galaxy-galaxy-$\kappa$ bispectrum in real data, in combination with the standard $2\times 2$-point data vector. We will use these measurements to carry out a first tentative cosmological analysis combining projected power spectra and bispectra, quantifying the ability of the bispectrum to reduce the final parameter uncertainties, and to test the robustness of cosmological constraints, through the consistency between different data combinations.
@@
The paper is structured as follows. Section 2 presents the FSB estimator, and the extension we introduce here to the case of multi-field cross-correlations. It also describes the model we will use to interpret our measurements, and our parameter inference pipeline. The galaxy samples and CMB lensing data we use in the analysis are described in Section 3. Our results, including an assessment of the gggggg and $gg\kappa$ bispectrum detections, the complementarity of different data combinations, and the cosmological constraints derived from the $gg$, $g\kappa$ and gggggg measurements, are presented in Section 4. Finally, Section 5 summarises our results and discusses potential avenues for future exploration.
@@ 2.1.1 The estimator
[…]
@@
[…]
@@
[…]
@@ 2.1.3 FSB covariance
    Pushing the FSB reinterpretation of the bispectrum as a power spectrum between $a$ and $a_{L}^{2}$, we can use existing techniques (García-García et al., 2019) to obtain a first analytical approximation of the covariance, by assuming that both $a$ and $a_{L}^{2}$ are Gaussian fields. Although this assumption is in general incorrect, since $a_{L}^{2}$ cannot be Gaussian even if $a$ is (which it generally is not), H25 shows it recovers all purely diagonal terms – both disconnected and connected – contributing to the covariance. This is not entirely surprising, as the “Gaussian” power spectrum covariance is known to dominate the statistical uncertainties even for the highly non-Gaussian fields commonly encountered in large-scale structure (Li et al., 2019).
@@ 2.2 Multi-field FSB
[…]
This measurement requires very few changes to the estimator: our original expression for the single-field FSB (Eq. 2) becomes
@@ 2.3 Maps and power spectrum measurements
  • To determine the detectability of the projected bispectrum, we will measure the FSB in four filters below the $2N_{\rm side}$ threshold. These four consecutive top-hat filters each have width $\Delta\ell=126$, partitioning the multipole range $\ell\in[2,\,506]$.
@@
This must be done for the filters and bandpower window functions corresponding to any field affected by the PWF (i.e. the galaxy overdensity, but not $\kappa$ in our case). Power spectra should be likewise corrected by the PWF where relevant. Note that any pure shot noise contribution (namely the power spectrum shot noise, $N_{\ell}^{gg}$, and the gggggg FSB noise term in Eq. 33) should not be corrected for the PWF, as white noise spectra remain constant regardless of the pixelization scheme [^4] This can be understood in terms of a purely Poisson process, in which the measurements on different pixels remain statistically independent regardless of the pixel size..
@@ 2.4.1 Galaxy clustering and CMB lensing
We will use a perturbative expansion in Eulerian space, up to second order, to connect the galaxy and matter overdensities (Fry and Gaztanaga, 1993; McDonald and Roy, 2009). Specifically, in Fourier space this reads
@@ 2.4.2 Power spectrum modelling
where $k_{\ell}(\chi)\equiv(\ell+1/2)/\chi$ is the 3D wavenumber associated with the angular scale $\ell$, and $q_{\alpha}(\chi)$ is the radial kernel of the $\alpha$ field. This relation is accurate in the Limber approximation (Limber, 1953; LoVerde and Afshordi, 2008), which holds for the wide radial kernels used in this work.
@@
where $P^{mm}(k,z)$ is the matter power spectrum. We also account for Redshift-Space Distortions (RSDs) in our modelling of the power spectra by adding a RSD term, linear in the matter density field, to the linear bias expansion.
@@ 2.4.3 Bispectrum Modelling
In this work we will focus on the galaxy-galaxy-galaxy and galaxy-galaxy-convergence bispectra (gggggg and $gg\kappa$ hereafter), and hence must model the 3D gggggg and ggmggm bispectra. We do not include the convergence-convergence-galaxy ($\kappa\kappa g$) bispectrum in this work, which might require a more complex modelling of the lensing reconstruction noise and is expected to have a lower signal than the gggggg and $gg\kappa$ bispectra. As in the case of the power spectrum, we will do this at lowest order in perturbation theory (i.e. tree level). Using the perturbative bias expansion introduced in Section 2.4.1, the corresponding bispectra are given by (see Schmittfull et al., 2015; Desjacques et al., 2018; Cagliari et al., 2025):
@@
$P_{k}$ denotes the linear matter power spectrum. The cyclical permutations for the gggggg bispectrum are, explicitly, $\{\mathbf{k},\mathbf{k}^{\prime}\}$, $\{\mathbf{k}^{\prime},\mathbf{k}^{\prime\prime}\}$, and $\{\mathbf{k}^{\prime\prime},\mathbf{k}\}$. In turn, the permutations in the $gg\kappa$ case are the same for the linear term, but only $\{\mathbf{k}^{\prime},\mathbf{k}^{\prime\prime}\}$ and $\{\mathbf{k}^{\prime\prime},\mathbf{k}\}$ contribute for the quadratic and tidal terms. Once computed, the 3D bispectrum $B^{ggg}(\mathbf{k},\mathbf{k}^{\prime},\mathbf{k}^{\prime\prime})$ ($B^{ggm}$) can be inserted into Eq. 2.4.3 to recover the projected bispectrum $b^{ggg}_{\ell\ell^{\prime}\ell^{\prime\prime}}$ ($b^{gg\kappa}_{\ell\ell^{\prime}\ell^{\prime\prime}}$) [^5] Note that we use the notation $b^{gg\kappa}$ (as opposed to $b^{ggm}$) to highlight the fact that we are using the CMB convergence $\kappa$ as tracer for the matter on the line of sight, $m$ – this relationship is encoded in the projection kernel $q_{\kappa}(\chi)$ (Eq. 16)..